OpenStudy (anonymous):

A wall clock is 12 inches in diameter. The clock has twelve equally spaced numbers. What is the distance around the edge of the clock from 8 to 12? π 4π 3π 2π

OpenStudy (anonymous):

@Hero

OpenStudy (ivettef365):

C = d. pir d=12 C = 12pi, so divide 12/12 x 4

OpenStudy (anonymous):

so 4

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

ok thanks

hero (hero):

This is a proportion problem dealing with distance vs degree measure. So we'll need to use the following formula: \[\frac{\text{distance from 8 - 12}}{\text{circumference of the circle}} = \frac{\text{degree measure from 8 - 12}}{360}\]

hero (hero):

The distance from 8 - 12 is what we need to find so we let that be x. The degree measure from 8 - 12 represents 1/4 of 360 or 45 degrees so. The circumference of the circle is 2pi*r or d*pi = 12pi \[\frac{x}{12\pi} = \frac{45}{360}\] Solve for x.

hero (hero):

@iFlexzatious

OpenStudy (anonymous):

thanks its 4

hero (hero):

Negative. It's not 4

hero (hero):

Solve the proportion above for x.

OpenStudy (anonymous):

I don't get I get 4pi

hero (hero):

It's not \(4 \pi\). Sorry to disapoint you.

hero (hero):

@ivettef365, why did you tell him \(4\pi\) was correct when it isn't?

OpenStudy (anonymous):

so then can u break it down for me more plz

hero (hero):

@iFlexzatious, do you know how to solve the proportion I posted above?

OpenStudy (anonymous):

no I don't understand that

hero (hero):

You don't understand how to solve this for x: \[\frac{x}{12\pi}= \frac{45}{360}\]

OpenStudy (anonymous):

cross mutiply

hero (hero):

45/360 reduces to 1/4 so: \[\frac{x}{12\pi} = \frac{1}{4}\]

hero (hero):

There, I just made it even easier for you.

OpenStudy (anonymous):

4x=12pi

OpenStudy (anonymous):

x=3

hero (hero):

\[x = 3\pi\]

OpenStudy (anonymous):

3pi

OpenStudy (anonymous):

thanks a lot can u come check this one out plz

OpenStudy (ivettef365):

still don't understand because from 8 - 12 is 1/3 of the Circumference C = 12 pi and 1/3 = 4

OpenStudy (anonymous):

45/360 reduces to 1/4 so:

hero (hero):

Actually, @ivettef365 is right.

hero (hero):

the proportion is \[\frac{x}{12\pi} = \frac{120}{360}\]

OpenStudy (anonymous):

cross multiply 4*x=4x 1*12=12pi therefore 4x=12pi 4/12pi=3pi @ivettef365

hero (hero):

\[\frac{x}{12\pi} = \frac{1}{3}\]

OpenStudy (anonymous):

3x=12pi x=4pi

hero (hero):

Which becomes \(3x = 12\pi\). \(x = 4\pi\)

OpenStudy (anonymous):

whats the right answer here